# A Structured Framework for Evaluating and Enhancing Interpretive Capabilities of Multimodal LLMs in Culturally Situated Tasks

### v1

@@ B.7 Guo Xi (Guo Xi)
    • “In real landscapes of rivers and valleys, observe them from afar to capture theirshi (shi - overall configuration/momentum), and observe them up close to capture theirzhi (zhi - substance/texture).”
@@
B.8 Dr. Aris Thorne (Dr. Aris Thorne)
@@ C.3 Structured Commentary Prompt Design
  • Brushwork and Technique: Analyze the use of lines (such as thickness, speed, turns, strength), the variations in ink tones (dense, light, wet, dry), texture strokes (texture strokes), moss dots (moss dots), coloring, and other specific painting techniques and their effects.
@@
  • Artistic Conception and Emotion (yijing): Elaborate on the overall atmosphere, aesthetic taste, and artistic style conveyed by the painting, as well as the emotional resonance or philosophical reflections it might evoke in the viewer.
@@ C.5.1 Evaluative Stance Labels
  • Historical Research (Historical Research)
  • Aesthetic Appreciation (Aesthetic Appreciation)
  • Socio-cultural Interpretation (Socio-cultural Interpretation)
  • Comparative Analysis ()
  • Theoretical Construction ()
  • Critical Inquiry (zhi)
  • High Praise ()
  • Objective Description ()
  • Mild Criticism ()
  • Strong Negation ()
@@ C.5.2 Core Focal Point Labels
  • Use of Color ()
  • Brushwork Technique ()
  • Texture Strokes (Chunfa) (texture strokes)
  • Line Quality (zhi)
  • Ink Application ()
  • Layout and Structure ()
  • Spatial Representation ()
  • Artistic Conception (yijing)
  • Emotional Expression ()
  • Subject Matter ()
  • Genre ()
  • Symbolism (yi)
  • Historical Context ()
  • Artist Biography (sheng)
  • Style/School ()
  • Technique Inheritance & Innovation ()
  • Cross-cultural Influence ()
@@ C.5.3 Argumentative Quality Labels
  • Profound Insight ()
  • Strong Argumentation ()
  • Clear Logic ()
  • Detailed Analysis ()
  • Classical Citations ()
  • Objective Viewpoint ()
  • Superficial Treatment ()
  • Overly General Content ()
  • Lacks Examples ()
  • Logical Gaps ()
  • Subjective/Biased View ()
@@ C.5.4 Derived Analytical Dimensions
  • Comprehensive Analyst Score ()
  • Historically Focused Critic Score ()
  • Technique & Style Focused Critic Score ()
  • Theory & Comparison Focused Critic Score ()
  • General Descriptive Profile Score ()
@@
  • Stylistic Analysis ()
  • Cross-cultural Comparison ()
  • Theoretical Construction ()
  • Overall Coherence Score ()
@@ C.6.1 Specialized Profile Criteria (Micro-Level)
  • (Technique & Style Focused):
@@
  • (Theory & Comparison Focused):
@@ C.6.2 General Descriptive Profile Criteria
  • (General Descriptive Profile):
@@ D.1 Detailed Persona Capability Scores
google_gemini-2.5pro Brother Thomas (Brother Thomas) -0.2 0.5 0.1 +6
@@
google_gemini-2.5pro Guo Xi (Guo Xi) -0.1 -0.1 0.2 +-7
google_gemini-2.5pro John Ruskin (John Ruskin) -0.2 0.5 0.2 +1
google_gemini-2.5pro Mama Zola (Mama Zola) -0.3 -0.0 0.1 +-2
google_gemini-2.5pro Su Shi (Su Shi) 0.4 0.5 0.4 +6
google_gemini-2.5pro Okakura Kakuz (Okakura Kakuzo) 0.1 0.3 0.1 +6
meta-llama_Llama-4-Scout-17B-16E-Instruct Brother Thomas (Brother Thomas) -0.1 0.1 -0.2 +6
@@
meta-llama_Llama-4-Scout-17B-16E-Instruct Guo Xi (Guo Xi) -0.3 -0.0 -0.4 +-3
meta-llama_Llama-4-Scout-17B-16E-Instruct John Ruskin (John Ruskin) 0.1 0.3 0.4 +0
meta-llama_Llama-4-Scout-17B-16E-Instruct Mama Zola (Mama Zola) -0.1 0.4 0.1 +2
meta-llama_Llama-4-Scout-17B-16E-Instruct Su Shi (Su Shi) -0.2 0.2 0.2 +-2
meta-llama_Llama-3.1-8B-Instruct Brother Thomas (Brother Thomas) -0.2 -0.2 -0.0 +0
@@
meta-llama_Llama-3.1-8B-Instruct Guo Xi (Guo Xi) 0.0 -0.9 -0.3 +-11
meta-llama_Llama-3.1-8B-Instruct John Ruskin (John Ruskin) -0.3 0.1 0.2 +-6
meta-llama_Llama-3.1-8B-Instruct Mama Zola (Mama Zola) -0.5 -0.4 -0.1 +-15
meta-llama_Llama-3.1-8B-Instruct Su Shi (Su Shi) 0.4 0.7 0.7 +10
Qwen-2.5-VL-7B Brother Thomas (Brother Thomas) 0.6 1.6 1.4 +19
@@
Qwen-2.5-VL-7B Guo Xi (Guo Xi) 0.5 1.2 1.0 +12
Qwen-2.5-VL-7B John Ruskin (John Ruskin) 0.7 1.7 1.3 +24
Qwen-2.5-VL-7B Mama Zola (Mama Zola) 0.9 2.4 2.1 +22
Qwen-2.5-VL-7B Su Shi (Su Shi) 0.8 1.5 1.5 +16
@@
  • Personas with Chinese cultural backgrounds (e.g., Mama Zola, Okakura Kakuz) generally scored higher in Chinese Painting Understanding and Chinese Language Usage.
@@ D.3 Supplementary Quantitative Data Tables
gemini-2.5pro model baseline -8.245 -7.489 august_(basic).txt
gemini-2.5pro model baseline -0.607 -15.201 august_(with_Dong_Qichang).txt
gemini-2.5pro model baseline -2.392 -1.717 august_(with_Dr_Evelyn_Reed).txt
gemini-2.5pro model baseline -12.369 -5.803 august_(with_Li_Ruoyun).txt
gemini-2.5pro model baseline -7.852 -6.419 august_(with_Marcus_Fabius).txt
human_expert human ground_truth 3.451 -0.876 JosephLevenson....txt
@@ Appendix E Knowledge Base Content
  • Chinese Landscape Painting Concepts ():
@@
    • Core Concept (): The core of Chinese landscape painting is "spirit resonance" (qi yun sheng dong), the foremost principle of Xie He’s "Six Canons", referring to the vitality, spirit, and verve presented in a work, emphasizing the unity of inner spirit and outer expression. Another core concept is "artistic conception" (yi jing), which is the emotion, atmosphere, and profound meaning conveyed by the painting beyond the objects themselves, pursuing an artistic effect of fused (emotion/scene) and jing (milieu/boundary), inspiring contemplation. Landscape painting also embodies the idea of "harmony between man and nature" (tian ren he yi), entrusting philosophical thoughts and emotions through the depiction of nature.
@@
    • Main Features (): The main features of Chinese landscape painting are: 1. Subject Matter: Primarily natural mountains and rivers, forests, clouds, and water, often imbued with literati sentiments such as reclusion and spiritual refreshment. 2. Brush and Ink (Bi Mo): Utilizes a brush, ink, and Xuan paper, emphasizing the "bone method in brushwork" (gu fa yong bi), shaping the texture of objects and expressing emotions through variations in the strength of lines and the density, wetness, and dryness of ink (e.g., outlining, texturing, rubbing, dotting, dyeing). 3. Composition (Zhang Fa): Focuses on the interplay of void and solid, appropriate density, echoing openings and closings, and leaving blank spaces to create profound artistic conception and pictorial momentum, often using perspective methods like "level distance" (ping yuan), "high distance" (gao yuan), and "deep distance" (shen yuan). 4. Pursuit of Artistic Conception: Seeks not complete formal resemblance but rather spiritual likeness, emphasizing the integration of poetry, calligraphy, painting, and seals, and pursuing meaning beyond the painted image.
@@
    • Brief History (): Chinese landscape painting originated in the Wei, Jin, Southern and Northern Dynasties, and became an independent genre in the Sui and Tang Dynasties. The Five Dynasties to the Northern Song (907-1127) was its "great era", with numerous famous artists (e.g., Jing Hao, Guan Tong, Dong Yuan, Ju Ran, Li Cheng, Fan Kuan, Guo Xi), forming distinct northern and southern styles: northern landscapes were majestic, while southern water towns were gentle. The Southern Song period placed more emphasis on poetic meaning and personal emotional expression (e.g., Ma Yuan, Xia Gui). Literati painting rose in the Yuan Dynasty, emphasizing the interest of brush and ink and subjective expression (e.g., Zhao Mengfu, the Four Masters of Yuan). The Ming and Qing Dynasties saw further development and a divergence of schools based on inherited traditions, with court painting and literati painting coexisting.
  • Qing Court Painting ():
@@
    • Overview (): Qing Dynasty court painting was managed by the Imperial Household Department. During the Qianlong era, specialized institutions such as the Ruyi Guan (Palace Ateliers) and the Painting Academy Office were established. Painters were strictly managed, with systems for examination, ranking, rewards and punishments, and work review. It primarily served the imperial family, with functions including recording the appearance and life of emperors and empresses, documenting major state events and ceremonies (e.g., Southern Inspection Tours, battle scenes), decorating palaces and gardens, religious propaganda, and historical reference. Its development is divided into three periods: Shunzhi-Kangxi (initial phase), Yongzheng-Qianlong (peak, with a complete system and numerous famous artists), and post-Jiaqing (decline), synchronized with the rise and fall of national strength.
@@
    • Characteristics (): Qing Dynasty court painting covered a wide range of subjects, including portraits of emperors, empresses, and meritorious officials, ’scenes of pleasure’ (xingletu), major historical events (Southern Inspection Tours, wars, ceremonies), religious paintings, decorative landscapes and flower-and-bird paintings, and documentary-style depictions of tribute animals and plants. The overall style was meticulous, detailed, richly colored, and regal. The most prominent characteristic was the fusion of Chinese and Western styles: influenced by European missionary painters, it emphasized light and shadow, three-dimensionality, employed linear perspective ("xianfa hua"), and introduced oil painting and copperplate engraving. Simultaneously, traditional landscape ("the Four Wangs" school) and flower-and-bird (Yun Shouping’s school) painting styles also continued.
@@
    • Representative Figures (): Representative painters include: early figures such as Jiao Bingzhen, Leng Mei, Tang Dai; peak period Chinese painters like Chen Mei, Ding Guanpeng, Jin Tingbiao, Xu Yang, Yao Wenhan, Zhang Zongcang; European painters (excluding Lang Shining) such as Jean Denis Attiret (Wang Zhicheng), Ignatius Sickeltart (Ai Qimeng), etc. Additionally, there were court official painters like Dong Bangda, Jiang Tingxi, etc.
  • Giuseppe Castiglione ():
@@
    • Biography Summary (sheng): Giuseppe Castiglione (Lang Shining, 1688-1766), an Italian from Milan, was a Jesuit. He came to China in the 54th year of Kangxi (1715) and entered the court around the Kangxi-Yongzheng transition, serving the Kangxi, Yongzheng, and Qianlong emperors. His main activities included creating paintings, participating in the design of the Western-style buildings in the Old Summer Palace (Yuanmingyuan), teaching Western painting techniques, and assisting Nian Xiyao in writing ’Shi Xue’ (The Study of Vision). He was favored during the Qianlong era and was posthumously granted the title of Vice Minister.
@@
    • Artistic Style Overview (): In his early period, Lang Shining’s style was typically Western. Later, to adapt to the aesthetic tastes of the Chinese imperial family, he integrated Chinese painting techniques, forming a style that blended Chinese and Western elements. His paintings emphasized realism, focusing on light and shadow, perspective, and anatomical structure, but also adopted Chinese painting methods such as even lighting and a focus on line work. Although his style was praised by the court, it was not recognized by the literati painting school.
@@
    • Major Contributions (): He systematically introduced Western painting techniques such as oil painting and linear perspective (xianfa hua) to the Qing court and taught them, promoting the fusion of Chinese and Western art and forming a new look for Qing court painting. He assisted in the completion of ’Shi Xue’ (The Study of Vision), advancing the spread of perspective studies. His documentary-style paintings are important historical materials.
@@
    • Representative Works Mention (): Besides the ’Twelve Months Paintings’, his representative works include ’One Hundred Horses’, ’Assembled Auspicious Objects’, ’Pine, Rock, and Auspicious Fungus’, ’Ayusi Attacking Bandits with a Spear’, ’Emperor Qianlong’s Spring Message of Peace’, etc. He also participated in creating large-scale documentary paintings such as ’Banquet in the Garden of Ten Thousand Trees’ and ’Equestrian Skills’.
  • Twelve Months Paintings (Twelve Months):
@@
    • Theme Content (): The ’Twelve Months Paintings’ is a series of 12 works on silk with colors, created by Lang Shining, depicting representative seasonal activities and life scenes in the Qing Dynasty court for each month of the year, such as viewing lanterns in the first month, dragon boat racing in the fifth month, and moon gazing in the eighth month, meticulously showcasing figures, costumes, architecture, and natural scenery.
@@
    • Artistic Significance (yi): This series is a mature representative work of Lang Shining’s style blending Chinese and Western elements, integrating Western perspective and light/shadow with traditional Chinese composition and aesthetics. It is not only a precious pictorial historical material for studying Qing Dynasty court life and culture but also an important testament to Sino-Western artistic exchange in the 18th century.
@@
    • Dataset Source Annotation (): The images for this research dataset are primarily sourced from the National Palace Museum (Taiwan) digital archives (600dpi, CC BY 4.0). Each painting has been annotated in three layers: visual elements, cultural symbols, and artistic techniques, to support AI evaluation and cultural-aesthetic analysis.

### v2

@@ B.7 Guo Xi (郭熙)
    • “In real landscapes of rivers and valleys, observe them from afar to capture their shi (势, overall configuration/momentum), and observe them up close to capture their zhi (质, substance/texture).”
@@
B.8 Dr. Aris Thorne (阿里斯·索恩博士)
@@ C.3 Structured Commentary Prompt Design
  • Brushwork and Technique: Analyze the use of lines (such as thickness, speed, turns, strength), the variations in ink tones (dense, light, wet, dry), texture strokes (皴法), moss dots (点苔), coloring, and other specific painting techniques and their effects.
@@
  • Artistic Conception and Emotion (意境): Elaborate on the overall atmosphere, aesthetic taste, and artistic style conveyed by the painting, as well as the emotional resonance or philosophical reflections it might evoke in the viewer.
@@ C.5.1 Evaluative Stance Labels
  • Historical Research (历史考证型)
  • Aesthetic Appreciation (美学鉴赏型)
  • Socio-cultural Interpretation (社会文化解读型)
  • Comparative Analysis (比较分析型)
  • Theoretical Construction (理论建构型)
  • Critical Inquiry (质疑与思辨型)
  • High Praise (高度赞扬与推崇)
  • Objective Description (客观中性描述)
  • Mild Criticism (温和批评与保留)
  • Strong Negation (强烈否定与驳斥)
@@ C.5.2 Core Focal Point Labels
  • Use of Color (色彩运用)
  • Brushwork Technique (笔法技巧)
  • Texture Strokes (皴法特点)
  • Line Quality (线条质量)
  • Ink Application (墨法变化)
  • Layout and Structure (布局与结构)
  • Spatial Representation (空间营造)
  • Artistic Conception (意境表达)
  • Emotional Expression (情感传递)
  • Subject Matter (主题内容)
  • Genre (题材选择)
  • Symbolism (象征意义)
  • Historical Context (历史背景)
  • Artist Biography (画家生平)
  • Style/School (风格流派)
  • Technique Inheritance & Innovation (技法传承与创新)
  • Cross-cultural Influence (跨文化影响)
@@ C.5.3 Argumentative Quality Labels
  • Profound Insight (见解深刻独到)
  • Strong Argumentation (论证充分有力)
  • Clear Logic (逻辑清晰严密)
  • Detailed Analysis (细节分析具体)
  • Classical Citations (引用经典佐证)
  • Objective Viewpoint (观点客观公允)
  • Superficial Treatment (论述流于表面)
  • Overly General Content (内容较为宽泛)
  • Lacks Examples (缺乏具体例证)
  • Logical Gaps (逻辑存在跳跃)
  • Subjective/Biased View (观点主观片面)
@@ C.5.4 Derived Analytical Dimensions
  • Comprehensive Analyst Score (博学通论型得分)
  • Historically Focused Critic Score (历史考据型得分)
  • Technique & Style Focused Critic Score (技艺风格型得分)
  • Theory & Comparison Focused Critic Score (理论比较型得分)
  • General Descriptive Profile Score (泛化描述型得分)
@@
  • Stylistic Analysis (风格分析)
  • Cross-cultural Comparison (跨文化比较)
  • Theoretical Construction (理论建构)
  • Overall Coherence Score (整体连贯性得分)
@@ C.6.1 Specialized Profile Criteria (Micro-Level)
  • 技艺风格型 (Technique & Style Focused):
@@
  • 理论比较型 (Theory & Comparison Focused):
@@ C.6.2 General Descriptive Profile Criteria
  • 泛化描述型 (General Descriptive Profile):
@@ D.1 Detailed Persona Capability Scores
google_gemini-2.5pro Brother Thomas (托马斯修士) -0.2 0.5 0.1 +6
@@
google_gemini-2.5pro Guo Xi (郭熙) -0.1 -0.1 0.2 +-7
google_gemini-2.5pro John Ruskin (约翰·罗斯金) -0.2 0.5 0.2 +1
google_gemini-2.5pro Mama Zola (佐拉妈妈) -0.3 -0.0 0.1 +-2
google_gemini-2.5pro Su Shi (苏轼) 0.4 0.5 0.4 +6
google_gemini-2.5pro Okakura Kakuzō (冈仓天心) 0.1 0.3 0.1 +6
meta-llama_Llama-4-Scout-17B-16E-Instruct Brother Thomas (托马斯修士) -0.1 0.1 -0.2 +6
@@
meta-llama_Llama-4-Scout-17B-16E-Instruct Guo Xi (郭熙) -0.3 -0.0 -0.4 +-3
meta-llama_Llama-4-Scout-17B-16E-Instruct John Ruskin (约翰·罗斯金) 0.1 0.3 0.4 +0
meta-llama_Llama-4-Scout-17B-16E-Instruct Mama Zola (佐拉妈妈) -0.1 0.4 0.1 +2
meta-llama_Llama-4-Scout-17B-16E-Instruct Su Shi (苏轼) -0.2 0.2 0.2 +-2
meta-llama_Llama-3.1-8B-Instruct Brother Thomas (托马斯修士) -0.2 -0.2 -0.0 +0
@@
meta-llama_Llama-3.1-8B-Instruct Guo Xi (郭熙) 0.0 -0.9 -0.3 +-11
meta-llama_Llama-3.1-8B-Instruct John Ruskin (约翰·罗斯金) -0.3 0.1 0.2 +-6
meta-llama_Llama-3.1-8B-Instruct Mama Zola (佐拉妈妈) -0.5 -0.4 -0.1 +-15
meta-llama_Llama-3.1-8B-Instruct Su Shi (苏轼) 0.4 0.7 0.7 +10
Qwen-2.5-VL-7B Brother Thomas (托马斯修士) 0.6 1.6 1.4 +19
@@
Qwen-2.5-VL-7B Guo Xi (郭熙) 0.5 1.2 1.0 +12
Qwen-2.5-VL-7B John Ruskin (约翰·罗斯金) 0.7 1.7 1.3 +24
Qwen-2.5-VL-7B Mama Zola (佐拉妈妈) 0.9 2.4 2.1 +22
Qwen-2.5-VL-7B Su Shi (苏轼) 0.8 1.5 1.5 +16
@@
  • Personas with Chinese cultural backgrounds (e.g., Mama Zola, Okakura Kakuzō) generally scored higher in Chinese Painting Understanding and Chinese Language Usage.
@@ D.3 Supplementary Quantitative Data Tables
gemini-2.5pro model baseline -8.245 -7.489 august_bayue(basic).txt
gemini-2.5pro model baseline -0.607 -15.201 august_bayue(with_Dong_Qichang).txt
gemini-2.5pro model baseline -2.392 -1.717 august_bayue(with_Dr_Evelyn_Reed).txt
gemini-2.5pro model baseline -12.369 -5.803 august_bayue(with_Li_Ruoyun).txt
gemini-2.5pro model baseline -7.852 -6.419 august_bayue(with_Marcus_Fabius).txt
human_expert human ground_truth 3.451 -0.876 Levenson_Joseph...Chinese_early_painting_political_personal_factors.txt
@@ Appendix E Knowledge Base Content
  • Chinese Landscape Painting Concepts (中国山水画概念):
@@
    • Core Concept (核心理念): The core of Chinese landscape painting is “spirit resonance” (qi yun sheng dong), the foremost principle of Xie He’s “Six Canons”, referring to the vitality, spirit, and verve presented in a work, emphasizing the unity of inner spirit and outer expression. Another core concept is “artistic conception” (yi jing), which is the emotion, atmosphere, and profound meaning conveyed by the painting beyond the objects themselves, pursuing an artistic effect of fused 情景 (emotion/scene) and 境 (milieu/boundary), inspiring contemplation. Landscape painting also embodies the idea of “harmony between man and nature” (tian ren he yi), entrusting philosophical thoughts and emotions through the depiction of nature.
@@
    • Main Features (主要特点): The main features of Chinese landscape painting are: 1. Subject Matter: Primarily natural mountains and rivers, forests, clouds, and water, often imbued with literati sentiments such as reclusion and spiritual refreshment. 2. Brush and Ink (bi mo): Utilizes a brush, ink, and Xuan paper, emphasizing the “bone method in brushwork” (gu fa yong bi), shaping the texture of objects and expressing emotions through variations in the strength of lines and the density, wetness, and dryness of ink (e.g., outlining, texturing, rubbing, dotting, dyeing). 3. Composition (zhang fa): Focuses on the interplay of void and solid, appropriate density, echoing openings and closings, and leaving blank spaces to create profound artistic conception and pictorial momentum, often using perspective methods like “level distance” (ping yuan), “high distance” (gao yuan), and “deep distance” (shen yuan). 4. Pursuit of Artistic Conception: Seeks not complete formal resemblance but rather spiritual likeness, emphasizing the integration of poetry, calligraphy, painting, and seals, and pursuing meaning beyond the painted image.
@@
    • Brief History (简史): Chinese landscape painting originated in the Wei, Jin, Southern and Northern Dynasties, and became an independent genre in the Sui and Tang Dynasties. The Five Dynasties to the Northern Song (907-1127) was its "great era", with numerous famous artists (e.g., Jing Hao, Guan Tong, Dong Yuan, Ju Ran, Li Cheng, Fan Kuan, Guo Xi), forming distinct northern and southern styles: northern landscapes were majestic, while southern water towns were gentle. The Southern Song period placed more emphasis on poetic meaning and personal emotional expression (e.g., Ma Yuan, Xia Gui). Literati painting rose in the Yuan Dynasty, emphasizing the interest of brush and ink and subjective expression (e.g., Zhao Mengfu, the Four Masters of Yuan). The Ming and Qing Dynasties saw further development and a divergence of schools based on inherited traditions, with court painting and literati painting coexisting.
  • Qing Court Painting (清代宫廷绘画):
@@
    • Overview (概述): Qing Dynasty court painting was managed by the Imperial Household Department. During the Qianlong era, specialized institutions such as the Ruyi Guan (Palace Ateliers) and the Painting Academy Office were established. Painters were strictly managed, with systems for examination, ranking, rewards and punishments, and work review. It primarily served the imperial family, with functions including recording the appearance and life of emperors and empresses, documenting major state events and ceremonies (e.g., Southern Inspection Tours, battle scenes), decorating palaces and gardens, religious propaganda, and historical reference. Its development is divided into three periods: Shunzhi-Kangxi (initial phase), Yongzheng-Qianlong (peak, with a complete system and numerous famous artists), and post-Jiaqing (decline), synchronized with the rise and fall of national strength.
@@
    • Characteristics (特点): Qing Dynasty court painting covered a wide range of subjects, including portraits of emperors, empresses, and meritorious officials, “scenes of pleasure” (xing le tu), major historical events (Southern Inspection Tours, wars, ceremonies), religious paintings, decorative landscapes and flower-and-bird paintings, and documentary-style depictions of tribute animals and plants. The overall style was meticulous, detailed, richly colored, and regal. The most prominent characteristic was the fusion of Chinese and Western styles: influenced by European missionary painters, it emphasized light and shadow, three-dimensionality, employed linear perspective (xian fa hua), and introduced oil painting and copperplate engraving. Simultaneously, traditional landscape ("the Four Wangs" school) and flower-and-bird (Yun Shouping’s school) painting styles also continued.
@@
    • Representative Figures (代表人物): Representative painters include: early figures such as Jiao Bingzhen, Leng Mei, Tang Dai; peak period Chinese painters like Chen Mei, Ding Guanpeng, Jin Tingbiao, Xu Yang, Yao Wenhan, Zhang Zongcang; European painters (excluding Lang Shining) such as Jean Denis Attiret (Wang Zhicheng), Ignatius Sickeltart (Ai Qimeng), etc. Additionally, there were court official painters like Dong Bangda, Jiang Tingxi, etc.
  • Giuseppe Castiglione (郎世宁):
@@
    • Biography Summary (生平简介): Giuseppe Castiglione (Lang Shining, 1688-1766), an Italian from Milan, was a Jesuit. He came to China in the 54th year of Kangxi (1715) and entered the court around the Kangxi-Yongzheng transition, serving the Kangxi, Yongzheng, and Qianlong emperors. His main activities included creating paintings, participating in the design of the Western-style buildings in the Old Summer Palace (Yuanmingyuan), teaching Western painting techniques, and assisting Nian Xiyao in writing ‘Shi Xue’ (The Study of Vision). He was favored during the Qianlong era and was posthumously granted the title of Vice Minister.
@@
    • Artistic Style Overview (艺术风格概述): In his early period, Lang Shining’s style was typically Western. Later, to adapt to the aesthetic tastes of the Chinese imperial family, he integrated Chinese painting techniques, forming a style that blended Chinese and Western elements. His paintings emphasized realism, focusing on light and shadow, perspective, and anatomical structure, but also adopted Chinese painting methods such as even lighting and a focus on line work. Although his style was praised by the court, it was not recognized by the literati painting school.
@@
    • Major Contributions (主要贡献): He systematically introduced Western painting techniques such as oil painting and linear perspective (xian fa hua) to the Qing court and taught them, promoting the fusion of Chinese and Western art and forming a new look for Qing court painting. He assisted in the completion of ‘Shi Xue’ (The Study of Vision), advancing the spread of perspective studies. His documentary-style paintings are important historical materials.
@@
    • Representative Works Mention (代表作列举): Besides the ‘Twelve Months Paintings’, his representative works include ‘One Hundred Horses’, ‘Assembled Auspicious Objects’, ‘Pine, Rock, and Auspicious Fungus’, ‘Ayusi Attacking Bandits with a Spear’, ‘Emperor Qianlong’s Spring Message of Peace’, etc. He also participated in creating large-scale documentary paintings such as ‘Banquet in the Garden of Ten Thousand Trees’ and ‘Equestrian Skills’.
  • Twelve Months Paintings (十二月令图):
@@
    • Theme Content (主题内容): The ‘Twelve Months Paintings’ is a series of 12 works on silk with colors, created by Lang Shining, depicting representative seasonal activities and life scenes in the Qing Dynasty court for each month of the year, such as viewing lanterns in the first month, dragon boat racing in the fifth month, and moon gazing in the eighth month, meticulously showcasing figures, costumes, architecture, and natural scenery.
@@
    • Artistic Significance (艺术意义): This series is a mature representative work of Lang Shining’s style blending Chinese and Western elements, integrating Western perspective and light/shadow with traditional Chinese composition and aesthetics. It is not only a precious pictorial historical material for studying Qing Dynasty court life and culture but also an important testament to Sino-Western artistic exchange in the 18th century.
@@
    • Dataset Source Annotation (数据集来源与标注): The images for this research dataset are primarily sourced from the National Palace Museum (Taiwan) digital archives (600dpi, CC BY 4.0). Each painting has been annotated in three layers: visual elements, cultural symbols, and artistic techniques, to support AI evaluation and cultural-aesthetic analysis.